### v1

# Graph-based RAG Enhancement via Global Query Disambiguation and Dependency-Aware Reranking

Abstract
Contemporary graph-based retrieval-augmented generation (RAG) methods typically begin by extracting entities from user queries and then leverage pre-constructed knowledge graphs to retrieve related relationships and metadata. However, this pipeline’s exclusive reliance on entity-level extraction can lead to the misinterpretation or omission of latent yet critical information and relations. As a result, retrieved content may be irrelevant or contradictory, and essential knowledge may be excluded, exacerbating hallucination risks and degrading the fidelity of generated responses. To address these limitations, we introduce PankRAG, a framework that combines a globally aware, hierarchical query-resolution strategy with a novel dependency-aware reranking mechanism. PankRAG first constructs a multi-level resolution path that captures both parallel and sequential interdependencies within a query, guiding large language models (LLMs) through structured reasoning. It then applies its dependency-aware reranker to exploit the dependency structure among resolved sub-questions, enriching and validating retrieval results for subsequent sub-questions. Empirical evaluations demonstrate that PankRAG consistently outperforms state-of-the-art approaches across multiple benchmarks, underscoring its robustness and generalizability.
Graph-based RAG Enhancement via Global Query Disambiguation and Dependency-Aware Reranking
Ningyuan Li1, Junrui Liu1, Yi Shan1 , Minghui Huang1  and Tong Li1 1 College of Computer Science Beijing university of technology ningy@emails.bjut.edu.cn, liujunrui@emails.bjut.edu.cn,litong@bjut.edu.cn
1 Introduction
Large language models (LLMs) exhibit exceptional performance across diverse tasks by leveraging pre-training on massive text corpora to generate appropriate responses. While LLMs’ capabilities can be further enhanced through fine-tuning on downstream tasks, finite parameter capacity limits their ability to store comprehensive and evolving world knowledge. Retrieval-augmented generation (RAG)  Lewis et al. (2020); Guu et al. (2020); Gao et al. (2023); Ram et al. (2023); Fan et al. (2024); Izacard et al. (2023); Gao et al. (2024)has emerged as the predominant solution, owing to its straightforward implementation. It enables static models to access up-to-date external knowledge while preserving interpretability and provenance tracking.
Retrieval-augmented generation (RAG) methods dynamically integrate heterogeneous knowledge bases Sun et al. (2023); Edge et al. (2024); Guo et al. (2024); Ma et al. (2024) to enhance contextual reasoning and factual grounding. Among these, knowledge graphs  Chen et al. (2020); Jin et al. (2024) serve as structured frameworks that encode information through interconnected entities and relationships, explicitly modeling semantic dependencies to enable human-like understanding and multi-hop logical inference. Recently, many approaches have integrated RAG with knowledge graphs. For example, GraphRAG Edge et al. (2024) enhances private-text question answering by precomputing community summaries for clusters of related entities, dynamically synthesizing these summaries into contextually unified responses during query processing. LightRAG  Guo et al. (2024) employs a graph-text hybrid architecture that embeds structural relationships into document indexing and retrieval. Its hierarchical retrieval system combines graph-based relational reasoning with vector similarity matching, enabling efficient knowledge extraction while preserving semantic fidelity in multi-hop queries.
However, current graph-based RAG methods exhibit critical limitations of inaccuracies in the interpretation of critical information or relationships when handling specific and abstract complex queries. Specific complex queries (SCQ) demand entity-centric granularity, requiring precise retrieval of discrete entities and their relational contexts within structured knowledge systems, e.g., "In which direction would you travel to reach Southampton from the city connected to it by rail?". For specific queries, existing methods fail to capture latent information embedded in multi-hop relationships. For example, when GraphRAG’s local search retrieves content for the prior specific query, it returns extensive information about “Southampton,” most of which, beyond the city’s geographic location and its rail-connected neighbor, is irrelevant and may exacerbate LLM hallucinations. Abstract complex queries(ACQ) emphasize conceptual synthesis beyond entity-level precision, addressing domain-wide implications or systemic interactions, e.g., “How can theoretical principles from programming language theory be made accessible to new developers?”,(A detailed case study analyzing this example ACQ is presented in subsection A.3). For ACQ, methods often overlook the implicit context underlying conceptual abstractions. When LightRAG retrieves content for the prior ACQ, it prioritizes formal syntax definitions, while teaching tools, like step-by-step tutorials, error-driven case studies, and diverse code examples—provides richer contextual guidance, enabling the LLM to generate more in-depth responses. These limitations underscore the insufficiency of single-phase, query-as-is retrieval paradigms in addressing real-world SCQ and ACQ, challenging current graph-based RAG methods.
In this paper, we propose PankRAG, which resolves queries through a hierarchical resolution path with global awareness, augmented by a dependency-aware reranking mechanism. Specifically, our approach leverages the robust planning capabilities of LLMs to generate globally-aware (i.e., holistically planned rather than stepwise) problem-solving pathways, explicitly revealing parallel and progressive relationships among sub-questions. Next, we refine complex queries by integrating the implicit critical information resolved from prior sub-question resolutions. This dual-level process produces a hierarchical resolution pathway with global awareness, enhancing interpretability through contextual grounding. It enables LLMs to dynamically disambiguate and contextualize ambiguous or polysemous queries before final answer generation. During execution, our dependency-aware reranking mechanism augments and reorders retrieved content for sub-questions dependent on prior resolutions by integrating their resolved dependencies. This process refines the retrieved context by integrating resolved sub-question dependencies, thereby enhancing the alignment between retrieved contexts and dependency-driven sub-question answers challenging for LLMs to capture and resolve.
We evaluate PankRAG on four benchmark datasets for ACQ and two for SCQ, demonstrating that it consistently outperforms state-of-the-art baselines across six benchmarks. Compared to NaiveRAG, GraphRAG, and LightRAG, PankRAG achieves average win rates of 87.0%, 59.78% and 61.59% respectively in ACQ datasets, and achieves average metric improvements of 17.42%, 11.61%, 11.08% in SCQ datasets.
Our contributions to this work are as follows:
  • We propose a hierarchical query decomposition strategy that systematically plans globally-aware problem-solving pathways, enhancing LLMs’ capacity to capture and resolve implicit information and relationships, fostering alignment between query intent and generated responses.
  • We propose a dependency-aware reranking mechanism that augments and reorders retrieved content for the current sub-question by integrating resolved sub-question dependencies.
  • PankRAG demonstrates consistent superiority over state-of-the-art baselines, achieving comprehensive improvements on four ACQ benchmarks evaluated across six dimensions and two SCQ benchmarks assessed via four metrics.
Figure 1: The overall workflow of PankRAG. During Globally-Aware Planning, queries are decomposed into colored resolution pathways. In the bottom-up phase, two parallel sub-question sequences are generated: a blue sequence for interdependent sub-questions that require sequential processing, and an green independent sequence that can be resolved concurrently. In the top-down phase, an orange sequence rewrites complex queries, incorporating resolved sub-question dependencies to enhance clarity.
2 Releated Works
2.1 Retrieval-Augmented Generation Based On Chunks
RAG Fan et al. (2024); Gao et al. (2024) is widely used to address the timeliness of knowledge bases and hallucination issues Yao et al. (2023) in large language models (LLMs). Additionally, this approach is more cost-effective and easier to implement compared to methods like fine-tuning or retraining, significantly reducing computational and time resources. NaiveRAG  Gao et al. (2024) is a method that typically processes documents and queries by splitting them into smaller text chunks, converting these chunks into vector representations. The retrieved chunks, which are selected by comparing similarity with queries in vector space, are then used to generate responses through LLMs.
Numerous advanced RAG methods, such as MemoRAG Qian et al. (2024), RQ-RAG Chan et al. (2024), and others Asai et al. (2023); Lin et al. (2023), build upon the foundational framework of NaiveRAG, introducing specialized improvements. MemoRAG  Qian et al. (2024) is an advanced RAG framework empowered by global memory-augmented retrieval. It combines two specialized components: a lightweight system that generates draft answers as "global memory" to guide precise retrieval, and a powerful system that crafts the final answer using the retrieved information. RQ-RAG is a method designed to tackle ambiguous or complex queries. The model is equipped with explicit capabilities for query rewriting, decomposition, and disambiguation to enhance retrieval accuracy. Inspired by two methods, our proposed method integrates query decomposition and query rewriting principles, enhancing comprehension of hidden critical information or relationships within queries.
However, most text-based RAG methods utilize a flat structured database, failing to capture critical inter-chunk relationships, which constrains the quality of LLM-generated responses. While advanced RAG variants improve generation accuracy, their reliance on retrieving a limited subset of documents restricts comprehensive global information integration. It ultimately hinders their ability to address tasks requiring full-context understanding, such as ACQ.
2.2 Retrieval-Augmented Generation Based On KGs
Apart from a database of texts, a knowledge graph Jin et al. (2024); Chen et al. (2020), a structured framework organizing knowledge through interconnected entities and relationships, is increasingly employed as a valuable source for RAG methods. Unlike traditional RAG Gao et al. (2024), Graph-based RAG methods Sun et al. (2023); Ma et al. (2024)leverage graphs either as knowledge carriers or indexing tools to enable efficient retrieval. By modeling relationships between nodes, they facilitate the discovery of topic-centered or entity-centered knowledge pieces, ensuring thorough and contextually coherent retrieval. Furthermore, the graph structure contained dependencies or relationships between entities, which enhances both precision and scalability compared to flat databases.
Graph-based RAG methods differ in how they leverage the underlying graph data. GraphRAG  Edge et al. (2024) is designed to enhance question answering over private text corpora by using a knowledge graph. It pre-generates community summaries for clusters of related entities and dynamically aggregates them into coherent final responses during query resolution. LightRAG  Guo et al. (2024) is a novel framework that integrates graph structures into text indexing and retrieval. Employing a dual-level retrieval mechanism, it enhances comprehensive knowledge discovery and accelerates response generation without compromising contextual accuracy.
While prior research introduces knowledge graphs as a structured database, inaccuracies in the interpretation of hidden critical information or relationships within queries remain persistent challenges. Specifically, while knowledge graphs provide RAG systems with robust knowledge representation and enhanced retrieval performance, their practical effectiveness depends on accurately comprehending latent critical information and relationships within queries to identify precise retrieval targets. Our proposed framework is specifically designed to address this challenge. We also notice a concurrent work, CogGRAG Cheng et al. (2025) that leverages the principle of query decomposition. While they predominantly focus on problem decomposition strategy, our research prioritizes the acquisition of problem-solving pathways, with particular emphasis on the interplay between parallel interrelationships within these pathways and the impact of resolved sub-question dependencies on contextualizing information retrieved for the current sub-question.
3 Methodology
In this section, we detail our proposed framework, which comprises three core components. First, Graph Construction extracts entities and relationships from external knowledge sources to build a structured knowledge graph. Next, Globally-Aware Planning formulates a query resolution pathway by identifying latent critical information and relationships, improving LLMs’ comprehensiveness of complex queries. Finally, during Execution, the LLM processes sub-questions either concurrently or sequentially based on their interrelationships. After retrieving context for a sub-question, the dependency-aware reranking mechanism augments and validates the content when the current sub-question depends on a prior sub-question. The LLM then leverages this refined context to generate accurate responses. The overall workflow of FG-RAG is shown in Figure 1.
3.1 Graph Contruction
The source document corpus $D$ is segmented into $n$ text chunks $\{D_{i}\}^{n}_{i=1}$. As the Equation 1, for each chunk $D_{i}$, the LLM performs entity recognition and relationship extraction, abstracting structured semantic elements including entities $V_{i}$ and relations $E_{i}$ from unstructured text via prompting strategies. A knowledge graph $G$ is then constructed by instantiating nodes representing entities $\{V_{i}\}$ and edges representing relationships $\{E_{i}\}$ in Equation 2.
$(\mathcal{V}_{i},\mathcal{E}_{i})=\bigcup_{D_{i}\in\mathcal{D}}\textit{Extract}(D_{i})$ (1)
$\mathcal{G}=\left\{\mathcal{V}_{i},\mathcal{E}_{i}\right\}_{i=1}^{n}$ (2)
Next, Multiple communities $\{Com_{i}\}$ are aggregated into a higher-level community $H_{i}$ via Leiden clustering, and this process is iteratively applied to hierarchically construct a multi-layered community structure. Finally, multi-scale community summarization is executed by the LLM, which populates a summary template with various elements (nodes, edges, and associated claims).
$\mathcal{H}_{i}=\textit{Leiden}((\textit{Com}_{i})_{i=1}^{k})$ (3)
3.2 Globally-Aware Planning
Figure 2: Operation example of Globally-Aware Planning. The color assignments follow the configuration illustrated in the caption of Figure 1. [A1.1], [A1.2], and [A2.1] correspond to the resolved outputs of sub-questions SubQ[1.1], SubQ[1.2], and SubQ[2.1], respectively.
We propose a hierarchical query decomposition strategy that synergizes bottom-up relationship decomposition with top-down contextual rephrasing to address both implicit parallel or progress relationships and ambiguous or polysemous expressions in complex queries. Unlike conventional iterative approaches (e.g., Chain-of-Thought, CoT), which suffer from fixed iteration counts and myopic decomposition prone to local optima, our strategy dynamically adjusts the granularity and sequence of reasoning steps based on query complexity.
In the bottom-up phase, two crucial relationship types are detected. First, parallelizable task relationships are identified to enable concurrent resolution of mutually independent sub-questions, optimizing computational efficiency by reducing redundant resource contention and minimizing latency. Second, latent dependencies are systematically extracted to ensure prerequisite sub-questions are solved before dependent tasks.
Simultaneously, the top-down phase disambiguates polysemous expressions into logically coherent substructures, combining the resolved sub-question dependencies. This dual mechanism reduces the risk of deviation in reasoning paths and enhances computational efficiency by eliminating redundant iterations, achieving a balanced trade-off between precision and scalability. The detailed workflow is presented below, and a real-world query in Multihop-RAG dataset example is above in Figure 2.
  • Bottom-up phase: In the Parallel Relationship Detection step, parallel semantic relationships—such as independent sub-questions or co-occurring entities—are identified. If no parallelism is detected, a single resolution sequence is initialized. When parallelism exists, multiple concurrent sequences (SubQ1.x and SubQ2.x in Figure 2) are adaptively spawned to optimize concurrent execution. Concurrently, latent progress relationships are extracted from the query. Upon identifying distinct relationships within the query, sub-questions are generated and formalized into a directed acyclic graph (DAG), where nodes represent sub-questions and edges encode their relationships. Finally, topological sorting is applied to enforce prerequisite-resolution order, ensuring logical coherence during execution.
  • Top-down phase: For queries containing explicit ambiguities or polysemous expressions, a disambiguation sequence combining resolved sub-question dependencies (SubQ0.x in Figure 2) is invoked. In this phase, an LLM-driven rephrasing process converts ambiguous queries into logically partitioned formulations.
3.3 Execution
3.3.1 Retrieval
Following GraphRAG, we apply distinct retrieval strategies for different query types.
For SCQ, we use a local search approach that employs vector similarity metrics to identify nodes explicitly referenced in the query, then expands to neighboring nodes to retrieve related relationships and community context. This method ensures precise grounding in entity-centric evidence while maintaining low computational overhead.
For ACQ, we perform hierarchical community retrieval, dynamically selecting between leaf-level summaries for detailed information and higher-level summaries for thematic overviews. This multi-scale strategy balances breadth and depth to search for relevant information for complex queries.
Table 1: Performance across four datasets and six evaluation dimensions in terms of averaged win rates.
Agriculture Legal CS Mix
NaiveRAG PankRAG NaiveRAG PankRAG NaiveRAG PankRAG NaiveRAG PankRAG
Comprehensiveness 18.0% 82.0% 9.6% 90.4% 20.4% 79.6% 16.8% 83.2%
Diversity 19.6% 80.4% 7.2% 92.8% 27.6% 72.4% 8.8% 91.2%
logicality 2.8% 97.2% 8.8% 91.2% 7.6% 92.4% 17.2% 82.8%
relevance 6.0% 94.0% 11.2% 88.8% 12.8% 87.2% 19.2% 80.8%
coherence 2.8% 97.2% 8.4% 91.6% 6.8% 93.2% 16.8% 83.2%
Empowerment 12.0% 86.0% 9.6% 90.4% 16.4% 83.6% 11.6% 88.4%
GraphRAG PankRAG GraphRAG PankRAG GraphRAG PankRAG GraphRAG PankRAG
Comprehensiveness 40.8% 59.2% 44.4% 55.6% 48.4% 53.2% 44.8% 55.2%
Diversity 44.8% 55.2% 39.2% 60.8% 39.6% 60.4% 43.6% 56.4%
logicality 32.8% 67.2% 43.6% 56.4% 46.4% 53.6% 30.0% 70.0%
relevance 40.4% 59.6% 38.0% 62.0% 39.6% 60.4% 44.4% 55.6%
coherence 34.4% 65.6% 38.0% 62.0% 42.4% 57.6% 31.2% 68.8%
Empowerment 39.6% 60.4% 36.4% 63.6% 44.0% 56.0% 40.0% 60.0%
LightRAG PankRAG LightRAG PankRAG LightRAG PankRAG LightRAG PankRAG
Comprehensiveness 44.8% 55.2% 48.4% 53.2% 35.0% 65.0% 45.2% 54.8%
Diversity 41.6% 58.4% 40.8% 71.6% 28.0% 72.0% 37.2% 62.8%
logicality 16.4% 83.6% 38.0% 57.2% 41.2% 58.8% 42.8% 57.6%
relevance 35.2% 64.8% 38.0% 56.0% 43.6% 56.4% 43.6% 56.4%
coherence 16.8% 83.2% 38.0% 55.6% 42.0% 58.0% 40.0% 60.0%
Empowerment 44.8% 55.2% 45.2% 58.4% 34.0% 60.0% 36.0% 64.0%
3.3.2 Dependency-Aware Reranking Mechanism
The dependency-aware reranking mechanism is triggered when resolving a sub-question that depends on previously resolved sub-questions. It dynamically adjusts retrieval priorities by indirectly leveraging these resolved dependencies. Unlike conventional methods that directly incorporate resolved sub-question dependencies, risking cumulative hallucination errors across iterative reasoning steps, our mechanism mitigates such propagation by reinforcing contextually relevant content and filtering out incongruent evidence through semantic similarity validation. If no dependencies are present, the generation proceeds without reranking, preserving computational efficiency while maintaining robustness against cascading inaccuracies.
The reranking mechanism computes a composite score by weighted summation of two components: the intrinsic quality score $R_{i}$ of retrieved content and the semantic similarity $M_{i}$ between retrieved content and resolved dependent sub-question answers. The intrinsic quality score $R_{i}$ of retrieved content is generated and retained during the original retrieval phase Edge et al. (2024). The semantic similarity $M_{i}$ is computed as the cosine similarity between resolved sub-question dependencies $\{sub_{i}\}$ and the initially retrieved text chunks $\{ori_{i}\}$. Each chunk’s similarity score is updated and then linearly combined with its original retrieval quality score to optimize prioritization. The weight parameters $\alpha$ (quality) and $\beta$ (relevance) were selected by testing on the dataset(subsection 4.4), ensuring maximal retrieval precision while mitigating hallucination propagation. This process can be formally described as follows.
$\mathcal{M}_{i}=\textit{Cos}(\textit{ori}_{i}+\textit{sub}_{i})$ (4)
Combined $\text{Combined Score}=\alpha\times\mathcal{R}_{i}+\beta\times\mathcal{M}_{i}$ (5)
4 Experiment
4.1 Experimental Setup
Table 2: The QA performance across two datasets and four key metrics: answer faithfulness, context recall, and context precision.
Multihop-RAG MuSiQue Average
Method Answer Relevance Faithfulness Context Recall Context Precision Answer Relevance Faithfulness Context Recall Context Precision Answer Relevance Faithfulness Context Recall Context Precision
NaiveRAG 66.73% 73.99% 55.17% 54.00% 66.70% 62.50% 59.41% 58.21% 66.72% 66.70% 57.29% 56.11%
GraphRAG 69.56% 86.80% 62.06% 61.18% 68.94% 69.73% 61.57% 60.40% 69.25% 0.78.27% 61.82% 60.70%
LightRAG 72.40% 86.78% 65.18% 67.00% 67.92% 68.68% 60.99% 58.42% 68.66% 77.73% 63.08% 0.6271%
PankRAG 0.7479 0.9653 0.8789 0.8400 0.7062 0.8875 0.6499 0.6535 0.7271 0.9264 0.7644 0.7468
4.1.1 Datasets
We adopt the experimental framework of recent graph-based RAG studies and evaluate our method on six real-world datasets, encompassing both SCQ and ACQ. We use the Musique and MultiHop-RAG datasets for SCQ, which require reasoning across multiple passages, understanding implicit relationships, and synthesizing information. ACQ demands holistic curation, multi-perspective synthesis, and coherent organization of information at scale. We evaluate our method on the UltraDomin dataset Qin et al. (2024), which is drawn from 428 college-level textbooks across 18 domains, selecting the four domains used (Agriculture, Legal, Computer Science, and Mix), whose detailed information is as follows in LightRAG’s evaluation. We follow the standardized preprocessing pipelines from GraphRAG and LightRAG.
Figure 3: Impact of a pair of hyperparameters in PankRAG on the Mix(left) and Agriculture(right) dataset
4.1.2 Baselines
We compare PathRAG with the following state-of-the-art methods across all datasets:
  • NaiveRAG Gao et al. (2024): In this RAG method, text is divided into chunks, stored as embeddings in a vector database, and the most semantically similar chunks are retrieved during queries to enable efficient matching.
  • GraphRAG Edge et al. (2024): In this graph-based RAG method, they hierarchically organize knowledge through entity-level descriptions and clustered community summaries. For global queries, it retrieves community-focused knowledge and aggregates related entities to produce comprehensive responses.
  • LightRAG Guo et al. (2024): In this graph-based RAG method, graph structures are integrated with vector representations, and a dual-level retrieval paradigm is employed to enhance both the efficiency and relevance of retrieval.
4.1.3 Evaluation Metrics
  • SCQ Dataset In the SCQ dataset, we evaluate all models using the Ragas Es et al. (2024) framework across four critical metrics: answer relevance, faithfulness, context recall, and context precision.
  • ACQ Dataset In the absence of ground-truth answers, we adopt LLM-based evaluation protocols following prior work. Specifically, we leverage GPT-4o-mini to assess generated answers across six dimensions: Comprehensiveness, Diversity and Empowerment, adopted from LightRAG Guo et al. (2024), and three novel dimensions—Logicality, Relevance, and Coherence-introduced by recent LLM evaluation research Chen et al. (2025). Answers generated by all baselines and our method are pairwise compared using win-rate analysis, where a higher win rate reflects superior performance relative to alternatives. To mitigate presentation bias, answer order is systematically alternated, with final results reporting averaged win rates. How much detail does the answer provide to cover all aspects and details of the question?
4.2 Overall Performance on the ACQ Datasets
As demonstrated in Table 1, PathRAG consistently surpasses baseline methods across all evaluation metrics and datasets.
From the perspective of benchmark datasets, PankRAG achieves average win rates of 72.47% on the Agriculture dataset, 69.87% on the Legal dataset, 67.77% on the CS dataset, and 68.4% on the Mix dataset. These results demonstrate PankRAG’s generalizability in effectively facilitating LLMs to capture hidden critical information and relationships in ACQ across datasets of varying scales, thereby achieving robust and superior performance.
From the evaluation dimension perspective, PankRAG attains average win rates of 65.55% in comprehensiveness, 69.64% in diversity, 72.33% in logicality, 68.5% in relevance, 73.00% in coherence, and 68.86% in empowerment. This consistent outperformance across all dimensions, attributed to the globally-aware pathway resolution and dependency-aware reranking mechanisms, underscores PankRAG’s comprehensive advantages over baseline RAG methods.
Figure 4: Impact of a pair of hyperparameters in PankRAG on the Multihop-RAG dataset
4.3 Overall Performance on the SCQ Datasets
As demonstrated in LABEL:tab:specificq_res, PankRAG consistently achieves the best performance across all experimental settings.
Compared to NaiveRAG, GraphRAG, and LightRAG, PankRAG achieves average metric improvements of 17.42%, 11.61%, 11.08%. Averaged over multiple datasets, the results are as follows. PankRAG achieves approximately 5.99% higher than NaiveRAG, 3.46% higher than GraphRAG, and 4.05% higher than LightRAG in Answer Relevance; 25.94% higher than NaiveRAG, 14.37% higher than GraphRAG, and 14.91% higher than LightRAG in Faithfulness; 19.15% higher than NaiveRAG, 14.62% higher than GraphRAG, and 13.36% higher than LightRAG in Context Recall; and 18.57% higher than NaiveRAG, 13.98% higher than GraphRAG, and 11.97% higher than LightRAG in Context Precision. These results demonstrate PankRAG’s superiority over prior graph-based RAG methods in handling concrete complex problems.
Notably, GraphRAG and LightRAG show minimal improvements in Context Recall and Context Precision (4%-6% over NaiveRAG), indicating negligible enhancement in retrieved context quality. This aligns with the identified challenge that LLMs struggle to interpret hidden relationships or information in SCQ, leading to misaligned retrieval targets. While these methods enhance retrieval via graph structures, their limited context quality gains highlight unresolved limitations. In contrast, PankRAG’s significant improvements in Context Recall/Precision suggest effective mitigation of this issue, attributed to its globally-aware pathway resolution generation module and dependency-aware reranking mechanism. It enables LLMs to better capture hidden relationships and thereby improve context relevance. Concurrent enhancements in Faithfulness and Answer Relevance further indicate reduced hallucination risks and higher answer quality, underscoring PankRAG’s comprehensive advancements over existing graph-based RAG approaches.
Table 3: Ablation study on abstract queries on the global-aware pathway resolution generation module and dependency-aware reranking mechanism in PankRAG for Abstract Queries.
Agriculture Legal CS Mix
PankRAG - Pank PankRAG - Ank PankRAG - Pank PankRAG - Ank PankRAG - Pank PankRAG - Ank PankRAG - Pank PankRAG - Ank
Comprehensiveness 46.8% 53.2% 34.8% 55.2% 47.2% 52.8% 46.0% 54.0%
Diversity 42.3% 58.8% 38.0% 62.0% 36.2% 54.8% 33.6% 56.4%
logicality 44.4% 55.6% 44.0% 56.0% 44.8% 55.2% 45.2% 54.8%
relevance 45.2% 54.8% 44.8% 55.2% 46.4% 53.6% 46.0% 54.0%
coherence 46.8% 53.2% 43.8% 57.2% 46.4% 53.6% 45.2% 54.8%
Empowerment 56.0% 54.0% 41.2% 58.8% 46.0% 54.0% 43.6% 54.4%
PankRAG - Ank PankRAG PankRAG - Ank PankRAG PankRAG - Ank PankRAG PankRAG - Ank PankRAG
Comprehensiveness 41.6% 58.4% 46.0% 54.0% 46.0% 54.0% 39.2% 55.2%
Diversity 46.4% 53.6% 44.8% 55.2% 41.2% 58.8% 30.8% 56.4%
logicality 33.2% 66.8% 43.2% 56.8% 44.8% 55.2% 32.4% 68.8%
relevance 41.6% 58.4% 46.0% 54.0% 45.6% 54.4% 32.4% 56.4%
coherence 36.8% 63.2% 43.0% 57.0% 43.6% 56.4% 32.4% 60.0%
Empowerment 37.2% 62.8% 43.4% 55.6% 44.0% 56.0% 42.4% 58.8%
Table 4: Ablation study on abstract queries on the global-aware pathway resolution generation module and dependency-aware reranking mechanism in PankRAG for Specific Queries.
Multihop-RAG MuSiQue Average
Method Answer Relevance Faithfulness Context Recall Context Precision Answer Relevance Faithfulness Context Precision Context Recall Answer Relevance Faithfulness Context Recall Context Precision
PankRAG-Pank 67.96 88.21% 62.94 62.21 67.45% 75.16% 59.90% 58.73% 67.70% 81.69% 61.42% 60.47%
PankRAG-Ank 70.21% 94.45 84.31 67.65 69.02% 81.13% 62.27% 61.23% 69.61% 87.79% 73.29% 68.63%
PankRAG 0.7479 0.9653 0.8789 0.8400 0.7062 0.8875 0.6499 0.6535 0.7271 0.9264 0.7644 0.7468
4.4 Hyperparameter Analysis
The parameters $\alpha$ and $\beta$ are designed to optimize the original quality score by integrating a novel similarity metric, constrained by $\alpha$ + $\beta$ = 1 to ensure the normalized score range. Experimental results demonstrate that the optimal parameters for SCQ are $\alpha$ > 0.6 $\beta$ < 0.4, whereas for ACQ, $\alpha$ = 0.75 and $\beta$ = 0.25 , with both configurations maximizing retrieval performance.
In SCQ datasets, we evaluated retrieval performance across various ($\alpha$:$\beta$) ratios in Figure 4. As ($\alpha$:$\beta$) increases, key metrics—including answer relevance, context recall, and context precision—monotonically improve, peaking at ($\alpha$:$\beta$) = 0.6:0.4, before declining. Although faithfulness attains its maximum at ($\alpha$:$\beta$) = 0.4:0.6, its value at $0.6:0.4$ is only 0.0027 lower, indicating a negligible trade-off. Thus, $\alpha$ = 0.6 and $\beta$ = 0.4 represent the optimal balance. We indicate that when $\alpha$ > 0.6 $\beta$ < 0.4, the weight on sub-question answers is insufficient to adequately reinforce and validate retrieved context. Conversely, when $\alpha$ < 0.6 $\beta$ > 0.4, over-emphasis on sub-question answers introduces negative hallucinations that outweigh positive reinforcement, degrading performance. At $\alpha$ = 0.6 and $\beta$ = 0.4, these opposing effects are balanced, yielding the best overall retrieval results.
In ACQ datasets, We adjust a pair of hyperparameters on the Mix and Agriculture dataset, and then calculate the win rates compared with LightRAG, the best baseline. As shown in  Figure 3, parameter configuration of $\alpha$ = 0.75 and $\beta$ = 0.25 achieves the highest win rates across the majority of the six dimensions and secures the second-highest performance in all remaining cases. For ACQ, the optimal parameter configuration that maximizes retrieval performance is achieved when $\alpha$ = 0.75 and $\beta$ = 0.25.
4.5 Ablation Studies
In ACQ datasets, Since our proposed dependency-aware reranking mechanism (Ank) operates on top of the globally-aware pathway resolution (Path), we first evaluated the effectiveness of Path by measuring the win rates of PankRAG-Path versus PankRAG-Ank across six dimensions, as shown in the Table 3. Subsequently, to validate the synergistic enhancement of Ank on the Path-integrated graph RAG framework, we compared the win rates of PankRAG-Ank against PankRAG across the same six dimensions.
As shown in Table 3, PathRAG achieves an average win rate of 57.75% against PathRAG-Ank, PathRAG-Ank attains a 55.27% average win rate over PathRAG-Pank. These results demonstrate the necessity of the globally-aware pathway resolution and further validate the dependency-aware reranking mechanism’s critical role in optimizing RAG systems incorporating globally-aware pathway resolution.
In SCQ datasets, we separately evaluated PankRAG-Pank, PankRAG-Ank, and PankRAG across four metrics to validate the effectiveness of the globally-aware pathway resolution (Path) and the dependency-aware rerank (Ank), respectively.
As shown in the Table 4, PankRAG surpasses PankRAG-Ank by approximately 3.1% and PankRAG-Pank by 4.51% in Answer Relevance. For Faithfulness, PankRAG outperforms PankRAG-Ank by 4.85% and PankRAG-Pank by 10.96%. In Context Precision, PankRAG exceeds PankRAG-Ank by 3.35% and PankRAG-Pank by 15.22%, while in Context Recall, it surpasses PankRAG-Ank by 6.05% and PankRAG-Pank by 14.21%. Notably, PankRAG-Ank achieves significant improvements over PankRAG-Pank in Context Precision and Recall, indicating enhanced retrieval performance and underscoring the necessity of the globally-aware pathway resolution. Concurrently, PankRAG’s consistent superiority over PankRAG-Ank across all metrics highlights the critical role of the dependency-aware reranking mechanism.
5 Conclusion
In this paper, we propose PankRAG, a novel RAG framework designed to enhance LLMs’ capacity to accurately interpret and retrieve information for complex queries. PankRAG combines a globally aware pathway-resolution module with a dependency-aware reranking mechanism. By executing globally coherent reasoning pathways and augmenting retrieved content through dependency-aware reranking, our framework captures and resolves hidden relationships and information in both abstract and specific complex queries, significantly enhancing retrieval performance and response quality.
Limitations
While PankRAG enhances retrieval accuracy and comprehensiveness through globally-aware query resolution paths and dependency-aware reranking mechanisms, it faces two key limitations. First, resolving multiple sub-problems during retrieval necessitates additional LLM invocations, which improves quality of response but introduces higher computational overhead. Second, our evaluation follows the relative metrics of existing graph-based RAG methods, which may lack stability compared to absolute benchmarks. Future work will focus on establishing more reliable and absolute evaluation metrics tailored to graph-based RAG frameworks.
Appendix A Appendix
A.1 Datasets
A.1.1 ACQ
  • CS: This dataset focuses on core computer science domains, emphasizing data science and software engineering. It highlights practical implementations in machine learning (e.g., recommendation systems, classification algorithms) and big data processing (e.g., Spark-based real-time analytics).
  • Legal: This dataset centers on corporate legal frameworks, covering restructuring procedures, compliance protocols, legal agreements, and governance mechanisms, with applications in legal-financial sector contexts.
  • Agriculture: This dataset addresses agricultural practices, including beekeeping strategies, crop production optimization, disease prevention techniques, and farm management methodologies.
  • Mix: This dataset compiles literary classics across genres (essays, poetry, biographies), spanning philosophy, history, and cultural studies through interdisciplinary perspectives.
A.2 Evaluation
A.2.1 ACQ
LLM evaluate RAG responses according to following seven dimensions:
  • Comprehensiveness: How much detail does the answer provide to cover all aspects and details of the question?
  • Diversity: How varied and rich is the answer in providing different perspectives and insights on the question?
  • Logicality: How logically does the answer respond to all parts of the question?
  • Relevance: How relevant is the answer to the question, staying focused and addressing the intended topic or issue?
  • Coherence: How well does the answer maintain internal logical connections between its parts, ensuring a smooth and consistent structure?
  • Empowerment: How well does the answer help the reader understand and make informed judgments about the topic?
To mitigate ordering effects in comparative evaluations, we implement a systematic rotation of answer positions between the two methods under investigation. Each method’s output is alternately assigned to the primary position (Answer A) and the secondary position (Answer B) across evaluation trials. This dual-perspective counterbalancing enables the computation of bidirectional win rates, ensuring evaluation parity while controlling for positional bias inherent in sequential prompt structures.
A.2.2 SCQ
  • This metric quantifies the degree to which generated responses directly address the original query {r_q}, emphasizing completeness and specificity while penalizing redundant or tangential content. Automated evaluation involves generating {r_n} hypothetical questions from the answer using LLMs, embedding both original and generated questions (via text-embedding-ada-002), and computing their cosine similarity. The final score is formally defined as:
    $\text{AR}=\frac{1}{n}\sum_{i=1}^{n}\cos(\mathbf{e}_{q},\mathbf{e}_{q_{i}})$ (6)
    where {e_q} denotes the original question and {e_q_i} represent LLM-generated counterpart questions.
  • Faithfulness(F): This metric measures the factual grounding of answers within provided contexts, critical for mitigating hallucinations. The evaluation involves: 1. Decomposing answers into atomic claims from 2. Verifying each claim against through entailment checks. The faithfulness score is calculated as:
    $F=\frac{|\text{Verified Claims}|}{|\text{Total Claims}|}$ (7)
    where {_V} denotes verified claims and {_S} the total extracted statements.
  • Context Recall(CR): This metric evaluates the precision of retrieved contexts by quantifying the proportion of context sentences directly relevant to answering the query. Using LLM extraction, critical sentences {S_ext} through structured prompting. The context are identified from {C_q} recall score (CR) is derived as:
    $\text{CR}=\frac{|S_{\text{ext}}|}{|c(q)|}$ (8)
    where {c(q)} represents the total sentences in the context. represents the total sentences in the context. Low CR scores indicate excessive irrelevant content, highlighting retrieval inefficiency.
A.3 Case Study
To demonstrate the advantages of PankRAG over existing RAG systems, we conduct a comparative analysis between PankRAG and baseline methods using the ACQ task ("How can theoretical principles from programming language theory be made accessible to new developers?") from the CS dataset. Due to space constraints, we present only the case study of the ACQ task with PankRAG versus LightRAG. As shown in the table, with the help of a hierarchical resolution path with global awareness in Figure 5, PankRAG outperforms LightRAG across all six evaluation metrics, with detailed analyses evaluated from LLMs provided below.
LightRAG Response Several approaches can facilitate the understanding of these principles, making them more relatable and useful for beginners embarking on their coding journeys. 1. Educational Resources : […] 2. Hands-On Experience : […] 3. Use of Simplified Language and Analogies : […] 4. Community Building and Support : […] Conclusion: Integrating educational resources, hands-on experience, simplified explanations, supportive communities, new developers can better access the theoretical principles of programming language theory…
PankRAG Response Making Theoretical Principles Accessible to New Developers To facilitate a smoother transition for new developers into the complexities of programming language theory … A structured approach that incorporates simplifications, practical engagement, interactive learning resources, and collaborative support will enhance comprehension and retention. 1. Simplification of Concepts : […] 2. Practical Engagement : […] 3. Interactive Learning Resources : […] 4. Collaborative Learning and Support : […] Conclusion: In summary, to make theoretical principles from programming language theory accessible to new developers, a multifaceted approach that emphasizes simplification, practical engagement, interactive resources, and collaborative support…
LLM Evaluation:
  • Comprehensiveness: “Winner”: "Answer2"
    “Explanation”: Answer2(PankRAG) provides a detailed breakdown of strategies to make theoretical principles accessible, including simplification, practical engagement, interactive resources, and collaborative support. It covers various aspects of the topic thoroughly, while Answer1(LightRAG) …
  • Diversity: “Winner”: "Answer2"
    “Explanation”: Answer2(PankRAG) presents a wider range of strategies and methods, including mentorship programs and specific examples of interactive tools. It emphasizes diverse learning approaches, while Answer1(LightRAG) …
  • Logicality: “Winner”: "Answer2"
    “Explanation”: Answer2(PankRAG) follows a clear, structured approach that logically progresses from simplification to practical engagement and collaborative support. Each section builds on the previous one, creating a cohesive argument. Answer1(LightRAG) …
  • Relevance: “Winner”: "Answer2"
    “Explanation”: Both answers are relevant to the question, but Answer2(PankRAG) maintains a sharper focus on practical strategies that directly address how to make theoretical principles accessible. Answer1(LightRAG) …
  • Coherence: “Winner”: "Answer2"
    “Explanation”:Answer2(PankRAG) exhibits a high degree of coherence, with each section clearly connected to the main theme and to each other. The flow of ideas is smooth and logical. Answer1(LightRAG) …
  • Empowerment: “Winner”: "Answer2"
    “Explanation”: Answer2(PankRAG) empowers the reader by providing actionable strategies and emphasizing the importance of mentorship and collaborative learning. It encourages new developers to engage actively with the material. Answer1(LightRAG) …
SubQ 1-1: What are the key theoretical principles from programming language theory? SubQ 2-1: What are effective methods or resources for teaching programming language theory to beginners? SubQ 1-2: What challenges do new developers face in understanding these principles? SubQ 0-1: How can the key theoretical principles be simplified or presented in a way that is more approachable for new developers? SubQ 0-2: What additional support or mentoring could be beneficial for new developers learning these concepts? Query: How can theoretical principles from programming language theory be made accessible to new developers?
Figure 5: A hierarchical resolution path with global awareness of ACQ

### v2

@@
-Graph-based RAG Enhancement via Global Query Disambiguation and Dependency-Aware Reranking
+PankRAG: Enhancing Graph Retrieval via Globally Aware Query Resolution and Dependency-Aware Reranking Mechanism
 Abstract
-Contemporary graph-based retrieval-augmented generation (RAG) methods typically begin by extracting entities from user queries and then leverage pre-constructed knowledge graphs to retrieve related relationships and metadata.
-However, this pipeline’s exclusive reliance on entity-level extraction can lead to the misinterpretation or omission of latent yet critical information and relations.
-As a result, retrieved content may be irrelevant or contradictory, and essential knowledge may be excluded, exacerbating hallucination risks and degrading the fidelity of generated responses.
-To address these limitations, we introduce PankRAG, a framework that combines a globally aware, hierarchical query-resolution strategy with a novel dependency-aware reranking mechanism.
-PankRAG first constructs a multi-level resolution path that captures both parallel and sequential interdependencies within a query, guiding large language models (LLMs) through structured reasoning.
-It then applies its dependency-aware reranker to exploit the dependency structure among resolved sub-questions, enriching and validating retrieval results for subsequent sub-questions.
-Empirical evaluations demonstrate that PankRAG consistently outperforms state-of-the-art approaches across multiple benchmarks, underscoring its robustness and generalizability.
-Graph-based RAG Enhancement via Global Query Disambiguation and Dependency-Aware Reranking
-Ningyuan Li1, Junrui Liu1, Yi Shan1 , Minghui Huang1  and Tong Li1 1 College of Computer Science Beijing university of technology ningy@emails.bjut.edu.cn, liujunrui@emails.bjut.edu.cn,litong@bjut.edu.cn
+Recent graph-based RAG approaches leverage knowledge graphs by extracting entities from a query to fetch their associated relationships and metadata.
+However, relying solely on entity extraction often results in the misinterpretation or omission of latent critical information and relationships.
+This can lead to the retrieval of irrelevant or contradictory content, as well as the exclusion of essential information, thereby increasing hallucination risks and undermining the quality of generated responses.
+In this paper, we propose PankRAG, a framework designed to capture and resolve the latent relationships within complex queries that prior methods overlook.
+It achieves this through a synergistic combination of a globally-aware hierarchical resolution pathway and a dependency-aware reranking mechanism.
+PankRAG first generates a globally aware resolution pathway that captures parallel and progress relationships, guiding LLMs to resolve queries through a hierarchical reasoning path.
+Additionally, its dependency-aware reranking mechanism utilizes resolved sub-question dependencies to augment and validate the retrieved content of the current unresolved sub-question.
+Experimental results demonstrate that PankRAG consistently outperforms existing state-of-the-art methods, underscoring its generalizability.
+The code is available at:https://github.com/Ninggggy/PankRAG.
+Index Terms—  Question answering, Retrieval-augmented generation, Query decomposition.
 1 Introduction
-Large language models (LLMs) exhibit exceptional performance across diverse tasks by leveraging pre-training on massive text corpora to generate appropriate responses.
-While LLMs’ capabilities can be further enhanced through fine-tuning on downstream tasks, finite parameter capacity limits their ability to store comprehensive and evolving world knowledge.
-Retrieval-augmented generation (RAG)  Lewis et al.
-(2020); Guu et al.
-(2020); Gao et al.
-(2023); Ram et al.
-(2023); Fan et al.
-(2024); Izacard et al.
-(2023); Gao et al.
-(2024)has emerged as the predominant solution, owing to its straightforward implementation.
+Large language models (LLMs)  [17, 18]exhibit exceptional performance across diverse tasks by leveraging pre-training [11] on massive text corpora to generate appropriate responses.
+While LLMs’ capabilities can be further enhanced through fine-tuning [2] on downstream tasks, finite parameter capacity limits their ability to store comprehensive and evolving world knowledge.
+Retrieval-augmented generation (RAG)  [15, 9, 6, 20, 5, 12, 7, 14]has emerged as the predominant solution, owing to its straightforward implementation.
 It enables static models to access up-to-date external knowledge while preserving interpretability and provenance tracking.
-Retrieval-augmented generation (RAG) methods dynamically integrate heterogeneous knowledge bases Sun et al.
-(2023); Edge et al.
-(2024); Guo et al.
-(2024); Ma et al.
-(2024) to enhance contextual reasoning and factual grounding.
-Among these, knowledge graphs  Chen et al.
-(2020); Jin et al.
-(2024) serve as structured frameworks that encode information through interconnected entities and relationships, explicitly modeling semantic dependencies to enable human-like understanding and multi-hop logical inference.
-Recently, many approaches have integrated RAG with knowledge graphs.
-For example, GraphRAG Edge et al.
-(2024) enhances private-text question answering by precomputing community summaries for clusters of related entities, dynamically synthesizing these summaries into contextually unified responses during query processing.
-LightRAG  Guo et al.
-(2024) employs a graph-text hybrid architecture that embeds structural relationships into document indexing and retrieval.
-Its hierarchical retrieval system combines graph-based relational reasoning with vector similarity matching, enabling efficient knowledge extraction while preserving semantic fidelity in multi-hop queries.
-However, current graph-based RAG methods exhibit critical limitations of inaccuracies in the interpretation of critical information or relationships when handling specific and abstract complex queries.
-Specific complex queries (SCQ) demand entity-centric granularity, requiring precise retrieval of discrete entities and their relational contexts within structured knowledge systems, e.g., "In which direction would you travel to reach Southampton from the city connected to it by rail?".
-For specific queries, existing methods fail to capture latent information embedded in multi-hop relationships.
-For example, when GraphRAG’s local search retrieves content for the prior specific query, it returns extensive information about “Southampton,” most of which, beyond the city’s geographic location and its rail-connected neighbor, is irrelevant and may exacerbate LLM hallucinations.
-Abstract complex queries(ACQ) emphasize conceptual synthesis beyond entity-level precision, addressing domain-wide implications or systemic interactions, e.g., “How can theoretical principles from programming language theory be made accessible to new developers?”,(A detailed case study analyzing this example ACQ is presented in subsection A.3).
-For ACQ, methods often overlook the implicit context underlying conceptual abstractions.
-When LightRAG retrieves content for the prior ACQ, it prioritizes formal syntax definitions, while teaching tools, like step-by-step tutorials, error-driven case studies, and diverse code examples—provides richer contextual guidance, enabling the LLM to generate more in-depth responses.
-These limitations underscore the insufficiency of single-phase, query-as-is retrieval paradigms in addressing real-world SCQ and ACQ, challenging current graph-based RAG methods.
-In this paper, we propose PankRAG, which resolves queries through a hierarchical resolution path with global awareness, augmented by a dependency-aware reranking mechanism.
-Specifically, our approach leverages the robust planning capabilities of LLMs to generate globally-aware (i.e., holistically planned rather than stepwise) problem-solving pathways, explicitly revealing parallel and progressive relationships among sub-questions.
+RAG dynamically integrate heterogeneous knowledge bases [21, 3, 8, 16, 13] to enhance contextual reasoning and factual grounding.
+Among these, knowledge graphs  [1] serve as structured frameworks that encode information through interconnected entities and relationships, explicitly modeling semantic dependencies to enable human-like understanding and multi-hop logical inference.
+Recently, many approaches have integrated RAG with knowledge graphs to enhance factual grounding and multi-hop inference, with prominent examples including GraphRAG[3], LightRAG[8], HippoRAG[13], and G-Retriever[10].
+For instance, GraphRAG[3] enhances question answering by adaptively synthesizing precomputed summaries of entity communities.
+LightRAG[8] leverages a hierarchical retrieval system that fuses graph-based reasoning with vector matching to ensure efficient knowledge extraction and semantic fidelity for multi-hop queries.
+However, the single-phase retrieval approach of current graph-based RAG methods is ill-suited for deep reasoning, causing them to misinterpret or omit latent critical information and relationships when handling complex queries.
+For Specific Complex Queries (SCQ) [8] requiring precise multi-hop inference, they often retrieve broad, irrelevant information centered on a core entity instead of the specific relational facts, increasing the risk of LLMs’ hallucinations.
+Similarly, for Abstract Complex Queries (ACQ) [8] demanding conceptual synthesis, they tend to fetch superficial definitions while overlooking the rich, implicit context required for a nuanced response.
+In this paper, we propose PankRAG, which resolves complex queries through a hierarchical reasoning pathway with global awareness, augmented by a dependency-aware reranking mechanism.
+Specifically, our approach leverages the robust planning capabilities of LLMs to generate a globally-aware (i.e., holistically planned rather than stepwise) resolution pathway, explicitly revealing parallel and progress relationships among sub-questions.
 Next, we refine complex queries by integrating the implicit critical information resolved from prior sub-question resolutions.
-This dual-level process produces a hierarchical resolution pathway with global awareness, enhancing interpretability through contextual grounding.
+This dual-level process produces a hierarchical resolution pathway with global awareness, significantly enhancing retrieval performance and response quality.
 It enables LLMs to dynamically disambiguate and contextualize ambiguous or polysemous queries before final answer generation.
-During execution, our dependency-aware reranking mechanism augments and reorders retrieved content for sub-questions dependent on prior resolutions by integrating their resolved dependencies.
-This process refines the retrieved context by integrating resolved sub-question dependencies, thereby enhancing the alignment between retrieved contexts and dependency-driven sub-question answers challenging for LLMs to capture and resolve.
-We evaluate PankRAG on four benchmark datasets for ACQ and two for SCQ, demonstrating that it consistently outperforms state-of-the-art baselines across six benchmarks.
-Compared to NaiveRAG, GraphRAG, and LightRAG, PankRAG achieves average win rates of 87.0%, 59.78% and 61.59% respectively in ACQ datasets, and achieves average metric improvements of 17.42%, 11.61%, 11.08% in SCQ datasets.
+During execution, our dependency-aware reranking mechanism augments and validates retrieved content for sub-questions.
+This process refines the retrieved context by integrating resolved sub-question dependencies, thereby enhancing the alignment between retrieved contexts and resolved sub-question dependencies.
 Our contributions to this work are as follows:
-  • We propose a hierarchical query decomposition strategy that systematically plans globally-aware problem-solving pathways, enhancing LLMs’ capacity to capture and resolve implicit information and relationships, fostering alignment between query intent and generated responses.
-  • We propose a dependency-aware reranking mechanism that augments and reorders retrieved content for the current sub-question by integrating resolved sub-question dependencies.
-  • PankRAG demonstrates consistent superiority over state-of-the-art baselines, achieving comprehensive improvements on four ACQ benchmarks evaluated across six dimensions and two SCQ benchmarks assessed via four metrics.
-Figure 1: The overall workflow of PankRAG.
-During Globally-Aware Planning, queries are decomposed into colored resolution pathways.
-In the bottom-up phase, two parallel sub-question sequences are generated: a blue sequence for interdependent sub-questions that require sequential processing, and an green independent sequence that can be resolved concurrently.
-In the top-down phase, an orange sequence rewrites complex queries, incorporating resolved sub-question dependencies to enhance clarity.
-2 Releated Works
-2.1 Retrieval-Augmented Generation Based On Chunks
-RAG Fan et al.
-(2024); Gao et al.
-(2024) is widely used to address the timeliness of knowledge bases and hallucination issues Yao et al.
-(2023) in large language models (LLMs).
-Additionally, this approach is more cost-effective and easier to implement compared to methods like fine-tuning or retraining, significantly reducing computational and time resources.
-NaiveRAG  Gao et al.
-(2024) is a method that typically processes documents and queries by splitting them into smaller text chunks, converting these chunks into vector representations.
-The retrieved chunks, which are selected by comparing similarity with queries in vector space, are then used to generate responses through LLMs.
-Numerous advanced RAG methods, such as MemoRAG Qian et al.
-(2024), RQ-RAG Chan et al.
-(2024), and others Asai et al.
-(2023); Lin et al.
-(2023), build upon the foundational framework of NaiveRAG, introducing specialized improvements.
-MemoRAG  Qian et al.
-(2024) is an advanced RAG framework empowered by global memory-augmented retrieval.
-It combines two specialized components: a lightweight system that generates draft answers as "global memory" to guide precise retrieval, and a powerful system that crafts the final answer using the retrieved information.
-RQ-RAG is a method designed to tackle ambiguous or complex queries.
-The model is equipped with explicit capabilities for query rewriting, decomposition, and disambiguation to enhance retrieval accuracy.
-Inspired by two methods, our proposed method integrates query decomposition and query rewriting principles, enhancing comprehension of hidden critical information or relationships within queries.
-However, most text-based RAG methods utilize a flat structured database, failing to capture critical inter-chunk relationships, which constrains the quality of LLM-generated responses.
-While advanced RAG variants improve generation accuracy, their reliance on retrieving a limited subset of documents restricts comprehensive global information integration.
-It ultimately hinders their ability to address tasks requiring full-context understanding, such as ACQ.
-2.2 Retrieval-Augmented Generation Based On KGs
-Apart from a database of texts, a knowledge graph Jin et al.
-(2024); Chen et al.
-(2020), a structured framework organizing knowledge through interconnected entities and relationships, is increasingly employed as a valuable source for RAG methods.
-Unlike traditional RAG Gao et al.
-(2024), Graph-based RAG methods Sun et al.
-(2023); Ma et al.
-(2024)leverage graphs either as knowledge carriers or indexing tools to enable efficient retrieval.
-By modeling relationships between nodes, they facilitate the discovery of topic-centered or entity-centered knowledge pieces, ensuring thorough and contextually coherent retrieval.
-Furthermore, the graph structure contained dependencies or relationships between entities, which enhances both precision and scalability compared to flat databases.
-Graph-based RAG methods differ in how they leverage the underlying graph data.
-GraphRAG  Edge et al.
-(2024) is designed to enhance question answering over private text corpora by using a knowledge graph.
-It pre-generates community summaries for clusters of related entities and dynamically aggregates them into coherent final responses during query resolution.
-LightRAG  Guo et al.
-(2024) is a novel framework that integrates graph structures into text indexing and retrieval.
-Employing a dual-level retrieval mechanism, it enhances comprehensive knowledge discovery and accelerates response generation without compromising contextual accuracy.
-While prior research introduces knowledge graphs as a structured database, inaccuracies in the interpretation of hidden critical information or relationships within queries remain persistent challenges.
-Specifically, while knowledge graphs provide RAG systems with robust knowledge representation and enhanced retrieval performance, their practical effectiveness depends on accurately comprehending latent critical information and relationships within queries to identify precise retrieval targets.
-Our proposed framework is specifically designed to address this challenge.
-We also notice a concurrent work, CogGRAG Cheng et al.
-(2025) that leverages the principle of query decomposition.
-While they predominantly focus on problem decomposition strategy, our research prioritizes the acquisition of problem-solving pathways, with particular emphasis on the interplay between parallel interrelationships within these pathways and the impact of resolved sub-question dependencies on contextualizing information retrieved for the current sub-question.
+  • We propose a hierarchical query decomposition strategy that generates a globally-aware resolution pathway, enabling LLMs to capture and resolve latent information and relationships, ensuring a tight alignment between generated responses and query intent.
+  • We propose a dependency-aware reranking mechanism that augments and validates retrieved content for the current sub-question by integrating resolved sub-question dependencies.
+  • PankRAG demonstrates consistent superiority over state-of-the-art baselines, achieving comprehensive improvements on four ACQ datasets evaluated across six dimensions and two SCQ datasets assessed via four metrics.
+Fig. 1: The overall workflow of PankRAG.
 2 Methodology
-In this section, we detail our proposed framework, which comprises three core components.
-First, Graph Construction extracts entities and relationships from external knowledge sources to build a structured knowledge graph.
-Next, Globally-Aware Planning formulates a query resolution pathway by identifying latent critical information and relationships, improving LLMs’ comprehensiveness of complex queries.
-Finally, during Execution, the LLM processes sub-questions either concurrently or sequentially based on their interrelationships.
-After retrieving context for a sub-question, the dependency-aware reranking mechanism augments and validates the content when the current sub-question depends on a prior sub-question.
-The LLM then leverages this refined context to generate accurate responses.
-The overall workflow of FG-RAG is shown in Figure 1.
 2.1 Graph Contruction
 The source document corpus $D$ is segmented into $n$ text chunks $\{D_{i}\}^{n}_{i=1}$.
-As the Equation 1, for each chunk $D_{i}$, the LLM performs entity recognition and relationship extraction, abstracting structured semantic elements including entities $V_{i}$ and relations $E_{i}$ from unstructured text via prompting strategies.
+As the Equation 1, for each chunk $D_{i}$, the LLM performs entity recognition and relationship extraction, abstracting entities $V_{i}$ and relations $E_{i}$ from unstructured text via prompting strategies.
 A knowledge graph $G$ is then constructed by instantiating nodes representing entities $\{V_{i}\}$ and edges representing relationships $\{E_{i}\}$ in Equation 2.
@@
 Next, Multiple communities $\{Com_{i}\}$ are aggregated into a higher-level community $H_{i}$ via Leiden clustering, and this process is iteratively applied to hierarchically construct a multi-layered community structure.
-Finally, multi-scale community summarization is executed by the LLM, which populates a summary template with various elements (nodes, edges, and associated claims).
+Finally, multi‑scale community summarization is executed by the LLM.
 $\mathcal{H}_{i}=\textit{Leiden}((\textit{Com}_{i})_{i=1}^{k})$ (3)
 2.2 Globally-Aware Planning
-Figure 2: Operation example of Globally-Aware Planning.
-The color assignments follow the configuration illustrated in the caption of Figure 1.
-[A1.1], [A1.2], and [A2.1] correspond to the resolved outputs of sub-questions SubQ[1.1], SubQ[1.2], and SubQ[2.1], respectively.
+Fig. 2: Operation example of Globally-Aware Planning. [A1.1], [A1.2], and [A2.1] correspond to the resolved outputs of SubQ[1.1], SubQ[1.2], and SubQ[2.1], respectively.
 We propose a hierarchical query decomposition strategy that synergizes bottom-up relationship decomposition with top-down contextual rephrasing to address both implicit parallel or progress relationships and ambiguous or polysemous expressions in complex queries.
-Unlike conventional iterative approaches (e.g., Chain-of-Thought, CoT), which suffer from fixed iteration counts and myopic decomposition prone to local optima, our strategy dynamically adjusts the granularity and sequence of reasoning steps based on query complexity.
-In the bottom-up phase, two crucial relationship types are detected.
-First, parallelizable task relationships are identified to enable concurrent resolution of mutually independent sub-questions, optimizing computational efficiency by reducing redundant resource contention and minimizing latency.
-Second, latent dependencies are systematically extracted to ensure prerequisite sub-questions are solved before dependent tasks.
-Simultaneously, the top-down phase disambiguates polysemous expressions into logically coherent substructures, combining the resolved sub-question dependencies.
-This dual mechanism reduces the risk of deviation in reasoning paths and enhances computational efficiency by eliminating redundant iterations, achieving a balanced trade-off between precision and scalability.
-The detailed workflow is presented below, and a real-world query in Multihop-RAG dataset example is above in Figure 2.
-  • Bottom-up phase: In the Parallel Relationship Detection step, parallel semantic relationships—such as independent sub-questions or co-occurring entities—are identified.
+Unlike conventional iterative approaches (e.g., Chain-of-Thought, CoT), which suffer from fixed iteration counts and myopic decomposition prone to local optima, our strategy dynamically adjusts the granularity and sequence of reasoning steps based on the query’s inherent structure.
+  • Bottom-up phase: In the Parallel Relationship Detection step, parallel semantic relationships—such as independent sub‑questions or co‑occurring entities—are identified to enable parallel execution, thereby optimizing computational efficiency by minimizing resource contention.
 If no parallelism is detected, a single resolution sequence is initialized.
@@
 Concurrently, latent progress relationships are extracted from the query.
-Upon identifying distinct relationships within the query, sub-questions are generated and formalized into a directed acyclic graph (DAG), where nodes represent sub-questions and edges encode their relationships.
-Finally, topological sorting is applied to enforce prerequisite-resolution order, ensuring logical coherence during execution.
+Upon identifying distinct relationships within the query, sub-questions are generated and formalized into a directed acyclic graph (DAG), where nodes represent sub‑questions and edges encode their relationships.
+Finally, topological sorting is applied to enforce prerequisite‑resolution order, ensuring logical coherence during execution.
   • Top-down phase: For queries containing explicit ambiguities or polysemous expressions, a disambiguation sequence combining resolved sub-question dependencies (SubQ0.x in Figure 2) is invoked.
 In this phase, an LLM-driven rephrasing process converts ambiguous queries into logically partitioned formulations.
+This dual mechanism reduces the risk of deviation in reasoning paths and enhances computational efficiency by eliminating redundant iterations, achieving a balanced trade‑off between precision and scalability.
+The detailed workflow is above inFigure 1, and a real-world query in the Multihop-RAG [22] dataset example is presented in Figure 2.
 2.3 Execution
-3.3.1 Retrieval
-Following GraphRAG, we apply distinct retrieval strategies for different query types.
-For SCQ, we use a local search approach that employs vector similarity metrics to identify nodes explicitly referenced in the query, then expands to neighboring nodes to retrieve related relationships and community context.
-This method ensures precise grounding in entity-centric evidence while maintaining low computational overhead.
-For ACQ, we perform hierarchical community retrieval, dynamically selecting between leaf-level summaries for detailed information and higher-level summaries for thematic overviews.
-This multi-scale strategy balances breadth and depth to search for relevant information for complex queries.
-Table 1: Performance across four datasets and six evaluation dimensions in terms of averaged win rates.
+Following GraphRAG[3], we tailor our retrieval strategy to the query type.
+For SCQs, we employ a local search that targets query-referenced nodes and expands to their neighbors, ensuring precise entity-centric grounding.
+For ACQs, we utilize hierarchical community retrieval, dynamically selecting between detailed leaf-level and thematic high-level summaries to balance retrieval breadth and depth.
+Table 1: Overall Performance on the ACQ Datasets.
 Agriculture Legal CS Mix
@@
 Empowerment 44.8% 55.2% 45.2% 58.4% 34.0% 60.0% 36.0% 64.0%
+HippoRAG PankRAG HippoRAG PankRAG HippoRAG PankRAG HippoRAG PankRAG
+Comprehensiveness 45.6% 54.4% 48.0% 52.0% 42.8% 57.2% 46.8% 53.2%
+Diversity 42.8% 57.2% 43.2% 56.8% 36.0% 74.0% 44.0% 56.0%
+logicality 31.2% 68.8% 43.6% 56.4% 42.8% 57.2% 44.4% 55.6%
+relevance 40.4% 59.6% 41.2% 58.8% 47.2% 52.8% 44.8% 55.2%
+coherence 32.4% 67.6% 42.4% 57.6% 39.6% 60.4% 45.2% 54.8%
+Empowerment 46.0% 54.0% 46.8% 53.2% 40.8% 59.2% 41.2% 58.8%
 2.3.1 Dependency-Aware Reranking Mechanism
-The dependency-aware reranking mechanism is triggered when resolving a sub-question that depends on previously resolved sub-questions.
+The dependency‑aware reranking mechanism is triggered when resolving a sub‑question that depends on previously resolved sub‑questions.
 It dynamically adjusts retrieval priorities by indirectly leveraging these resolved dependencies.
@@
 The reranking mechanism computes a composite score by weighted summation of two components: the intrinsic quality score $R_{i}$ of retrieved content and the semantic similarity $M_{i}$ between retrieved content and resolved dependent sub-question answers.
-The intrinsic quality score $R_{i}$ of retrieved content is generated and retained during the original retrieval phase Edge et al.
-(2024).
+$\mathcal{M}_{i}=\textit{Cos}(\textit{ori}_{i}+\textit{sub}_{i})$ (4)
+The intrinsic quality score $R_{i}$ of retrieved content is generated and retained during the original retrieval phase [3].
 The semantic similarity $M_{i}$ is computed as the cosine similarity between resolved sub-question dependencies $\{sub_{i}\}$ and the initially retrieved text chunks $\{ori_{i}\}$.
@@
 This process can be formally described as follows.
-$\mathcal{M}_{i}=\textit{Cos}(\textit{ori}_{i}+\textit{sub}_{i})$ (4)
-Combined $\text{Combined Score}=\alpha\times\mathcal{R}_{i}+\beta\times\mathcal{M}_{i}$ (5)
+$\text{Combined Score}=\alpha\times\mathcal{R}_{i}+\beta\times\mathcal{M}_{i}$ (5)
+2.3.2 Generation
+$Ans=\textit{LLM}(Q,(D,K))$ (6)
+The generation process integrates resolved dependencies $D$ and dependency-enhanced reranking knowledge $K$ into a consolidated context, which is utilized by the LLM to generate reliable responses.
 3 Experiment
 3.1 Experimental Setup
-Table 2: The QA performance across two datasets and four key metrics: answer faithfulness, context recall, and context precision.
+Table 2: Overall Performance on the SCQ Datasets
 Multihop-RAG MuSiQue Average
@@
 NaiveRAG 66.73% 73.99% 55.17% 54.00% 66.70% 62.50% 59.41% 58.21% 66.72% 66.70% 57.29% 56.11%
-GraphRAG 69.56% 86.80% 62.06% 61.18% 68.94% 69.73% 61.57% 60.40% 69.25% 0.78.27% 61.82% 60.70%
-LightRAG 72.40% 86.78% 65.18% 67.00% 67.92% 68.68% 60.99% 58.42% 68.66% 77.73% 63.08% 0.6271%
-PankRAG 0.7479 0.9653 0.8789 0.8400 0.7062 0.8875 0.6499 0.6535 0.7271 0.9264 0.7644 0.7468
-4.1.1 Datasets
-We adopt the experimental framework of recent graph-based RAG studies and evaluate our method on six real-world datasets, encompassing both SCQ and ACQ.
-We use the Musique and MultiHop-RAG datasets for SCQ, which require reasoning across multiple passages, understanding implicit relationships, and synthesizing information.
-ACQ demands holistic curation, multi-perspective synthesis, and coherent organization of information at scale.
-We evaluate our method on the UltraDomin dataset Qin et al.
-(2024), which is drawn from 428 college-level textbooks across 18 domains, selecting the four domains used (Agriculture, Legal, Computer Science, and Mix), whose detailed information is as follows in LightRAG’s evaluation.
-We follow the standardized preprocessing pipelines from GraphRAG and LightRAG.
-Figure 3: Impact of a pair of hyperparameters in PankRAG on the Mix(left) and Agriculture(right) dataset
-4.1.2 Baselines
-We compare PathRAG with the following state-of-the-art methods across all datasets:
-  • NaiveRAG Gao et al.
-(2024): In this RAG method, text is divided into chunks, stored as embeddings in a vector database, and the most semantically similar chunks are retrieved during queries to enable efficient matching.
-  • GraphRAG Edge et al.
-(2024): In this graph-based RAG method, they hierarchically organize knowledge through entity-level descriptions and clustered community summaries.
-For global queries, it retrieves community-focused knowledge and aggregates related entities to produce comprehensive responses.
-  • LightRAG Guo et al.
-(2024): In this graph-based RAG method, graph structures are integrated with vector representations, and a dual-level retrieval paradigm is employed to enhance both the efficiency and relevance of retrieval.
-4.1.3 Evaluation Metrics
-  • SCQ Dataset In the SCQ dataset, we evaluate all models using the Ragas Es et al.
-(2024) framework across four critical metrics: answer relevance, faithfulness, context recall, and context precision.
-  • ACQ Dataset In the absence of ground-truth answers, we adopt LLM-based evaluation protocols following prior work.
-Specifically, we leverage GPT-4o-mini to assess generated answers across six dimensions: Comprehensiveness, Diversity and Empowerment, adopted from LightRAG Guo et al.
-(2024), and three novel dimensions—Logicality, Relevance, and Coherence-introduced by recent LLM evaluation research Chen et al.
-(2025).
-Answers generated by all baselines and our method are pairwise compared using win-rate analysis, where a higher win rate reflects superior performance relative to alternatives.
-To mitigate presentation bias, answer order is systematically alternated, with final results reporting averaged win rates.
-How much detail does the answer provide to cover all aspects and details of the question?
+GraphRAG 69.56% 86.80% 62.06% 61.18% 68.94% 69.73% 61.57% 60.40% 69.25% 78.27% 61.82% 60.70%
+LightRAG 72.40% 86.78% 65.18% 67.00% 67.92% 68.68% 60.99% 58.42% 68.66% 77.73% 63.08% 62.71%
+HippoRAG 72.81% 89.84% 73.69% 72.35% 69.18% 75.68% 62.46% 62.19% 70.98% 82.76% 68.01% 67.27%
+PankRAG 74.79% 96.53% 87.89% 84.00% 70.62% 88.75% 64.99% 65.35% 72.71% 92.64% 76.44% 74.68%
+  • Datasets We follow the standardized preprocessing pipelines from GraphRAG[3] and LightRAG[8] and evaluate our method on six real‑world datasets, encompassing both SCQ and ACQ.
+We use the Musique [23] and MultiHop-RAG [22] datasets for SCQ.
+For ACQ, We evaluate our method on the UltraDomin dataset [19], selecting the four domains used (Agriculture, Legal, Computer Science, and Mix)
+  • Baselines We compare PankRAG with the three state-of-the-art methods, including NaiveRAG[20], GraphRAG[3], LightRAG[8] and HippoRAG[13] across all datasets.
+  • Evaluation Metrics On the SCQ datasets, we employ four metrics from the Ragas framework [4] to assess our method and the baselines: Answer Relevance, Faithfulness, Context Recall, and Context Precision.
+Lacking ground-truth for the ACQ dataset, we employ GPT-4o-mini as an automated judge.
+We conduct blinded, pairwise comparisons between Our evaluation involves calculating win rates for our method and baselines across six dimensions.
+To mitigate positional bias, the presentation order of answers is alternated, and we report the averaged results.
+We employed GPT-4o-mini for all core tasks.
+For text embeddings, we utilized the BAAI/bge-m3 model.
+Fig. 3: Impact of a pair of hyperparameters in PankRAG on the Mix(left) and Agriculture(right) dataset
 3.2 Overall Performance on the ACQ Datasets
-As demonstrated in Table 1, PathRAG consistently surpasses baseline methods across all evaluation metrics and datasets.
-From the perspective of benchmark datasets, PankRAG achieves average win rates of 72.47% on the Agriculture dataset, 69.87% on the Legal dataset, 67.77% on the CS dataset, and 68.4% on the Mix dataset.
-These results demonstrate PankRAG’s generalizability in effectively facilitating LLMs to capture hidden critical information and relationships in ACQ across datasets of varying scales, thereby achieving robust and superior performance.
-From the evaluation dimension perspective, PankRAG attains average win rates of 65.55% in comprehensiveness, 69.64% in diversity, 72.33% in logicality, 68.5% in relevance, 73.00% in coherence, and 68.86% in empowerment.
-This consistent outperformance across all dimensions, attributed to the globally-aware pathway resolution and dependency-aware reranking mechanisms, underscores PankRAG’s comprehensive advantages over baseline RAG methods.
-Figure 4: Impact of a pair of hyperparameters in PankRAG on the Multihop-RAG dataset
+As demonstrated in Table 1, PankRAG consistently surpasses baseline methods across all evaluation metrics and datasets.
+PankRAG demonstrates comprehensive superiority, achieving high average win rates across diverse datasets, ranging from 64% to 74%.
+This dominance extends across all evaluation dimensions, excelling particularly in Relevance (69.8%) and Empowerment (69.7%).
+This consistent outperformance is a direct result of its core mechanisms—globally-aware pathway resolution and dependency-aware reranking—validating its effectiveness on ACQ.
+Fig. 4: Impact of a pair of hyperparameters in PankRAG on the Multihop-RAG dataset[22]
+Table 3: Ablation study of PankRAG on ACQ datasets.
+Multihop-RAG MuSiQue Average
+Method Answer Relevance Faithfulness Context Recall Context Precision Answer Relevance Faithfulness Context Precision Context Recall Answer Relevance Faithfulness Context Recall Context Precision
+PankRAG-Pank 67.96% 88.21% 62.94 62.21 67.45% 75.16% 59.90% 58.73% 67.70% 81.69% 61.42% 60.47%
+PankRAG-Ank 70.21% 94.45 84.31 67.65 69.02% 81.13% 62.27% 61.23% 69.61% 87.79% 73.29% 68.63%
+PankRAG 74.79% 96.53% 87.89% 84.00% 70.62% 88.75% 64.99% 65.35% 72.71% 92.64% 76.44% 74.68%
 3.3 Overall Performance on the SCQ Datasets
-As demonstrated in LABEL:tab:specificq_res, PankRAG consistently achieves the best performance across all experimental settings.
-Compared to NaiveRAG, GraphRAG, and LightRAG, PankRAG achieves average metric improvements of 17.42%, 11.61%, 11.08%.
-Averaged over multiple datasets, the results are as follows.
-PankRAG achieves approximately 5.99% higher than NaiveRAG, 3.46% higher than GraphRAG, and 4.05% higher than LightRAG in Answer Relevance; 25.94% higher than NaiveRAG, 14.37% higher than GraphRAG, and 14.91% higher than LightRAG in Faithfulness; 19.15% higher than NaiveRAG, 14.62% higher than GraphRAG, and 13.36% higher than LightRAG in Context Recall; and 18.57% higher than NaiveRAG, 13.98% higher than GraphRAG, and 11.97% higher than LightRAG in Context Precision.
-These results demonstrate PankRAG’s superiority over prior graph-based RAG methods in handling concrete complex problems.
-Notably, GraphRAG and LightRAG show minimal improvements in Context Recall and Context Precision (4%-6% over NaiveRAG), indicating negligible enhancement in retrieved context quality.
+As demonstrated in Table 2, PankRAG consistently achieves the best performance across all experimental settings.
+PankRAG establishes its superiority by outperforming baselines by 11-17% on average.
+This advantage is driven by substantial gains across key metrics, including Faithfulness (10-26%), Context Precision(7-18%), Context Recall (8-19%), and Answer Relevance(2-6%), underscoring its enhanced efficacy in resolving SCQs.
+Notably, GraphRAG[3] and LightRAG[8] show minimal improvements in Context Recall and Context Precision (4%-6% over NaiveRAG), indicating negligible enhancement in retrieved context quality.
 This aligns with the identified challenge that LLMs struggle to interpret hidden relationships or information in SCQ, leading to misaligned retrieval targets.
-While these methods enhance retrieval via graph structures, their limited context quality gains highlight unresolved limitations.
-In contrast, PankRAG’s significant improvements in Context Recall/Precision suggest effective mitigation of this issue, attributed to its globally-aware pathway resolution generation module and dependency-aware reranking mechanism.
+In contrast, PankRAG’s significant improvements in Context Recall and Context Precision suggest effective mitigation of this issue, attributed to its globally-aware pathway resolution generation module and dependency-aware reranking mechanism.
 It enables LLMs to better capture hidden relationships and thereby improve context relevance.
-Concurrent enhancements in Faithfulness and Answer Relevance further indicate reduced hallucination risks and higher answer quality, underscoring PankRAG’s comprehensive advancements over existing graph-based RAG approaches.
-Table 3: Ablation study on abstract queries on the global-aware pathway resolution generation module and dependency-aware reranking mechanism in PankRAG for Abstract Queries.
+Concurrent enhancements in Faithfulness and Answer Relevance further indicate reduced hallucination risks and higher answer quality, underscoring PankRAG’s comprehensive superiority over existing graph-based RAG methods.
+Table 4: Ablation study of PankRAG on ACQ datasets.
 Agriculture Legal CS Mix
-PankRAG - Pank PankRAG - Ank PankRAG - Pank PankRAG - Ank PankRAG - Pank PankRAG - Ank PankRAG - Pank PankRAG - Ank
+w/o Pank w/o Ank w/o Pank w/o Ank w/o Pank w/o Ank w/o Pank w/o Ank
 Comprehensiveness 46.8% 53.2% 34.8% 55.2% 47.2% 52.8% 46.0% 54.0%
@@
 relevance 45.2% 54.8% 44.8% 55.2% 46.4% 53.6% 46.0% 54.0%
-coherence 46.8% 53.2% 43.8% 57.2% 46.4% 53.6% 45.2% 54.8%
+coherence 46.8% 53.2% 43.8% 7.2% 46.4% 53.6% 45.2% 54.8%
 Empowerment 56.0% 54.0% 41.2% 58.8% 46.0% 54.0% 43.6% 54.4%
-PankRAG - Ank PankRAG PankRAG - Ank PankRAG PankRAG - Ank PankRAG PankRAG - Ank PankRAG
+w/o Ank PankRAG w/o Ank PankRAG w/o Ank PankRAG w/o Ank PankRAG
 Comprehensiveness 41.6% 58.4% 46.0% 54.0% 46.0% 54.0% 39.2% 55.2%
@@
 Empowerment 37.2% 62.8% 43.4% 55.6% 44.0% 56.0% 42.4% 58.8%
-Table 4: Ablation study on abstract queries on the global-aware pathway resolution generation module and dependency-aware reranking mechanism in PankRAG for Specific Queries.
-Multihop-RAG MuSiQue Average
-Method Answer Relevance Faithfulness Context Recall Context Precision Answer Relevance Faithfulness Context Precision Context Recall Answer Relevance Faithfulness Context Recall Context Precision
-PankRAG-Pank 67.96 88.21% 62.94 62.21 67.45% 75.16% 59.90% 58.73% 67.70% 81.69% 61.42% 60.47%
-PankRAG-Ank 70.21% 94.45 84.31 67.65 69.02% 81.13% 62.27% 61.23% 69.61% 87.79% 73.29% 68.63%
-PankRAG 0.7479 0.9653 0.8789 0.8400 0.7062 0.8875 0.6499 0.6535 0.7271 0.9264 0.7644 0.7468
 3.4 Hyperparameter Analysis
-The parameters $\alpha$ and $\beta$ are designed to optimize the original quality score by integrating a novel similarity metric, constrained by $\alpha$ + $\beta$ = 1 to ensure the normalized score range.
-Experimental results demonstrate that the optimal parameters for SCQ are $\alpha$ > 0.6 $\beta$ < 0.4, whereas for ACQ, $\alpha$ = 0.75 and $\beta$ = 0.25 , with both configurations maximizing retrieval performance.
-In SCQ datasets, we evaluated retrieval performance across various ($\alpha$:$\beta$) ratios in Figure 4.
-As ($\alpha$:$\beta$) increases, key metrics—including answer relevance, context recall, and context precision—monotonically improve, peaking at ($\alpha$:$\beta$) = 0.6:0.4, before declining.
-Although faithfulness attains its maximum at ($\alpha$:$\beta$) = 0.4:0.6, its value at $0.6:0.4$ is only 0.0027 lower, indicating a negligible trade-off.
-Thus, $\alpha$ = 0.6 and $\beta$ = 0.4 represent the optimal balance.
-We indicate that when $\alpha$ > 0.6 $\beta$ < 0.4, the weight on sub-question answers is insufficient to adequately reinforce and validate retrieved context.
-Conversely, when $\alpha$ < 0.6 $\beta$ > 0.4, over-emphasis on sub-question answers introduces negative hallucinations that outweigh positive reinforcement, degrading performance.
-At $\alpha$ = 0.6 and $\beta$ = 0.4, these opposing effects are balanced, yielding the best overall retrieval results.
-In ACQ datasets, We adjust a pair of hyperparameters on the Mix and Agriculture dataset, and then calculate the win rates compared with LightRAG, the best baseline.
-As shown in  Figure 3, parameter configuration of $\alpha$ = 0.75 and $\beta$ = 0.25 achieves the highest win rates across the majority of the six dimensions and secures the second-highest performance in all remaining cases.
-For ACQ, the optimal parameter configuration that maximizes retrieval performance is achieved when $\alpha$ = 0.75 and $\beta$ = 0.25.
+The parameters $\alpha$ and $\beta$ (constrained by $\alpha$ + $\beta$ = 1 to ensure the normalized score range) balance the original quality score against a novel similarity metric derived from resolved dependencies.
+As shown in Figure 3 and Figure 4, Experimental results demonstrate that the optimal parameters for SCQ are $\alpha$ >0.6 $\beta$ <0.4, whereas for ACQ, $\alpha$ = 0.75 and $\beta$ = 0.25 , with both configurations maximizing retrieval performance.
+For SCQ, the optimal performance is achieved with $\alpha$ = 0.6 and $\beta$ = 0.4.
+This configuration represents a critical trade-off.
+A higher weight on $\alpha$ ($\alpha$ >0.6) results in insufficient reinforcement from sub-question answers, while a lower weight ($\alpha$ <0.6) leads to an over-emphasis that introduces negative hallucinations, degrading performance.
+The 0.6:0.4 ratio, therefore, strikes the ideal balance between leveraging sub-questions for context validation and mitigating the risk of performance degradation, yielding the best overall retrieval results.
+For ACQ, our experiments identified a different optimal configuration of $\alpha$ = 0.75 and $\beta$ = 0.25.
+This setting was determined as it consistently maximized retrieval performance and achieved the highest win rates against baseline models.
 3.5 Ablation Studies
-In ACQ datasets, Since our proposed dependency-aware reranking mechanism (Ank) operates on top of the globally-aware pathway resolution (Path), we first evaluated the effectiveness of Path by measuring the win rates of PankRAG-Path versus PankRAG-Ank across six dimensions, as shown in the Table 3.
-Subsequently, to validate the synergistic enhancement of Ank on the Path-integrated graph RAG framework, we compared the win rates of PankRAG-Ank against PankRAG across the same six dimensions.
-As shown in Table 3, PathRAG achieves an average win rate of 57.75% against PathRAG-Ank, PathRAG-Ank attains a 55.27% average win rate over PathRAG-Pank.
-These results demonstrate the necessity of the globally-aware pathway resolution and further validate the dependency-aware reranking mechanism’s critical role in optimizing RAG systems incorporating globally-aware pathway resolution.
-In SCQ datasets, we separately evaluated PankRAG-Pank, PankRAG-Ank, and PankRAG across four metrics to validate the effectiveness of the globally-aware pathway resolution (Path) and the dependency-aware rerank (Ank), respectively.
-As shown in the Table 4, PankRAG surpasses PankRAG-Ank by approximately 3.1% and PankRAG-Pank by 4.51% in Answer Relevance.
-For Faithfulness, PankRAG outperforms PankRAG-Ank by 4.85% and PankRAG-Pank by 10.96%.
-In Context Precision, PankRAG exceeds PankRAG-Ank by 3.35% and PankRAG-Pank by 15.22%, while in Context Recall, it surpasses PankRAG-Ank by 6.05% and PankRAG-Pank by 14.21%.
-Notably, PankRAG-Ank achieves significant improvements over PankRAG-Pank in Context Precision and Recall, indicating enhanced retrieval performance and underscoring the necessity of the globally-aware pathway resolution.
-Concurrently, PankRAG’s consistent superiority over PankRAG-Ank across all metrics highlights the critical role of the dependency-aware reranking mechanism.
+This section’s ablation studies validate the hierarchical contributions of our two core components: globally-aware pathway resolution (Path) and dependency-aware reranking (Ank).
+As shown in Table 3 and Table 4, the results demonstrate that the Path module is a foundational component crucial for enhancing retrieval quality.
+The model variant equipped with Path (PankRAG-Ank) achieves significant improvements in Context Precision and Context Recall over the variant without it (PankRAG-Pank).
+This finding underscores the necessity of the globally-aware pathway resolution for establishing a high-quality context.
+Furthermore, the Ank mechanism provides a critical, synergistic enhancement on top of the Path framework.
+The full PankRAG system, which incorporates the Ank module, shows consistent superiority over the PankRAG-Ank variant across all evaluated metrics.
+This highlights the critical role of the dependency-aware reranking mechanism in optimizing and refining the final results.
 4 Conclusion
 In this paper, we propose PankRAG, a novel RAG framework designed to enhance LLMs’ capacity to accurately interpret and retrieve information for complex queries.
-PankRAG combines a globally aware pathway-resolution module with a dependency-aware reranking mechanism.
-By executing globally coherent reasoning pathways and augmenting retrieved content through dependency-aware reranking, our framework captures and resolves hidden relationships and information in both abstract and specific complex queries, significantly enhancing retrieval performance and response quality.
-Limitations
-While PankRAG enhances retrieval accuracy and comprehensiveness through globally-aware query resolution paths and dependency-aware reranking mechanisms, it faces two key limitations.
-First, resolving multiple sub-problems during retrieval necessitates additional LLM invocations, which improves quality of response but introduces higher computational overhead.
-Second, our evaluation follows the relative metrics of existing graph-based RAG methods, which may lack stability compared to absolute benchmarks.
-Future work will focus on establishing more reliable and absolute evaluation metrics tailored to graph-based RAG frameworks.
-Appendix A Appendix
-A.1 Datasets
-A.1.1 ACQ
-  • CS: This dataset focuses on core computer science domains, emphasizing data science and software engineering.
-It highlights practical implementations in machine learning (e.g., recommendation systems, classification algorithms) and big data processing (e.g., Spark-based real-time analytics).
-  • Legal: This dataset centers on corporate legal frameworks, covering restructuring procedures, compliance protocols, legal agreements, and governance mechanisms, with applications in legal-financial sector contexts.
-  • Agriculture: This dataset addresses agricultural practices, including beekeeping strategies, crop production optimization, disease prevention techniques, and farm management methodologies.
-  • Mix: This dataset compiles literary classics across genres (essays, poetry, biographies), spanning philosophy, history, and cultural studies through interdisciplinary perspectives.
-A.2 Evaluation
-A.2.1 ACQ
-LLM evaluate RAG responses according to following seven dimensions:
-  • Comprehensiveness: How much detail does the answer provide to cover all aspects and details of the question?
-  • Diversity: How varied and rich is the answer in providing different perspectives and insights on the question?
-  • Logicality: How logically does the answer respond to all parts of the question?
-  • Relevance: How relevant is the answer to the question, staying focused and addressing the intended topic or issue?
-  • Coherence: How well does the answer maintain internal logical connections between its parts, ensuring a smooth and consistent structure?
-  • Empowerment: How well does the answer help the reader understand and make informed judgments about the topic?
-To mitigate ordering effects in comparative evaluations, we implement a systematic rotation of answer positions between the two methods under investigation.
-Each method’s output is alternately assigned to the primary position (Answer A) and the secondary position (Answer B) across evaluation trials.
-This dual-perspective counterbalancing enables the computation of bidirectional win rates, ensuring evaluation parity while controlling for positional bias inherent in sequential prompt structures.
-A.2.2 SCQ
-  • This metric quantifies the degree to which generated responses directly address the original query {r_q}, emphasizing completeness and specificity while penalizing redundant or tangential content.
-Automated evaluation involves generating {r_n} hypothetical questions from the answer using LLMs, embedding both original and generated questions (via text-embedding-ada-002), and computing their cosine similarity.
-The final score is formally defined as:
-    $\text{AR}=\frac{1}{n}\sum_{i=1}^{n}\cos(\mathbf{e}_{q},\mathbf{e}_{q_{i}})$ (6)
-    where {e_q} denotes the original question and {e_q_i} represent LLM-generated counterpart questions.
-  • Faithfulness(F): This metric measures the factual grounding of answers within provided contexts, critical for mitigating hallucinations.
-The evaluation involves: 1.
-Decomposing answers into atomic claims from 2.
-Verifying each claim against through entailment checks.
-The faithfulness score is calculated as:
-    $F=\frac{|\text{Verified Claims}|}{|\text{Total Claims}|}$ (7)
-    where {_V} denotes verified claims and {_S} the total extracted statements.
-  • Context Recall(CR): This metric evaluates the precision of retrieved contexts by quantifying the proportion of context sentences directly relevant to answering the query.
-Using LLM extraction, critical sentences {S_ext} through structured prompting.
-The context are identified from {C_q} recall score (CR) is derived as:
-    $\text{CR}=\frac{|S_{\text{ext}}|}{|c(q)|}$ (8)
-    where {c(q)} represents the total sentences in the context. represents the total sentences in the context. Low CR scores indicate excessive irrelevant content, highlighting retrieval inefficiency.
-A.3 Case Study
-To demonstrate the advantages of PankRAG over existing RAG systems, we conduct a comparative analysis between PankRAG and baseline methods using the ACQ task ("How can theoretical principles from programming language theory be made accessible to new developers?") from the CS dataset.
-Due to space constraints, we present only the case study of the ACQ task with PankRAG versus LightRAG.
-As shown in the table, with the help of a hierarchical resolution path with global awareness in Figure 5, PankRAG outperforms LightRAG across all six evaluation metrics, with detailed analyses evaluated from LLMs provided below.
-LightRAG Response Several approaches can facilitate the understanding of these principles, making them more relatable and useful for beginners embarking on their coding journeys.
-1.
-Educational Resources : […] 2.
-Hands-On Experience : […] 3.
-Use of Simplified Language and Analogies : […] 4.
-Community Building and Support : […] Conclusion: Integrating educational resources, hands-on experience, simplified explanations, supportive communities, new developers can better access the theoretical principles of programming language theory…
-PankRAG Response Making Theoretical Principles Accessible to New Developers To facilitate a smoother transition for new developers into the complexities of programming language theory … A structured approach that incorporates simplifications, practical engagement, interactive learning resources, and collaborative support will enhance comprehension and retention.
-1.
-Simplification of Concepts : […] 2.
-Practical Engagement : […] 3.
-Interactive Learning Resources : […] 4.
-Collaborative Learning and Support : […] Conclusion: In summary, to make theoretical principles from programming language theory accessible to new developers, a multifaceted approach that emphasizes simplification, practical engagement, interactive resources, and collaborative support…
-LLM Evaluation:
-  • Comprehensiveness: “Winner”: "Answer2"
-    “Explanation”: Answer2(PankRAG) provides a detailed breakdown of strategies to make theoretical principles accessible, including simplification, practical engagement, interactive resources, and collaborative support.
-It covers various aspects of the topic thoroughly, while Answer1(LightRAG) …
-  • Diversity: “Winner”: "Answer2"
-    “Explanation”: Answer2(PankRAG) presents a wider range of strategies and methods, including mentorship programs and specific examples of interactive tools.
-It emphasizes diverse learning approaches, while Answer1(LightRAG) …
-  • Logicality: “Winner”: "Answer2"
-    “Explanation”: Answer2(PankRAG) follows a clear, structured approach that logically progresses from simplification to practical engagement and collaborative support.
-Each section builds on the previous one, creating a cohesive argument.
-Answer1(LightRAG) …
-  • Relevance: “Winner”: "Answer2"
-    “Explanation”: Both answers are relevant to the question, but Answer2(PankRAG) maintains a sharper focus on practical strategies that directly address how to make theoretical principles accessible.
-Answer1(LightRAG) …
-  • Coherence: “Winner”: "Answer2"
-    “Explanation”:Answer2(PankRAG) exhibits a high degree of coherence, with each section clearly connected to the main theme and to each other.
-The flow of ideas is smooth and logical.
-Answer1(LightRAG) …
-  • Empowerment: “Winner”: "Answer2"
-    “Explanation”: Answer2(PankRAG) empowers the reader by providing actionable strategies and emphasizing the importance of mentorship and collaborative learning.
-It encourages new developers to engage actively with the material.
-Answer1(LightRAG) …
-SubQ 1-1: What are the key theoretical principles from programming language theory?
-SubQ 2-1: What are effective methods or resources for teaching programming language theory to beginners?
-SubQ 1-2: What challenges do new developers face in understanding these principles?
-SubQ 0-1: How can the key theoretical principles be simplified or presented in a way that is more approachable for new developers?
-SubQ 0-2: What additional support or mentoring could be beneficial for new developers learning these concepts?
-Query: How can theoretical principles from programming language theory be made accessible to new developers?
-Figure 5: A hierarchical resolution path with global awareness of ACQ
+PankRAG combines a globally aware pathway resolution module with a dependency‐aware reranking mechanism.
+By executing a globally coherent reasoning pathway and augmenting retrieved content through dependency‐aware reranking, our framework captures and resolves hidden relationships and information in both SCQ and ACQ, significantly enhancing retrieval performance and response quality.